# Learning Mean-Field Games through Mean-Field Actor-Critic Flow

### v1

@@ B.2 Convergence of the gap for value function
$\leq\int_{B_{R_{1}}}\frac{7}{3}\eta\rho_{0}(x)\,\mathrm{d}x+\int_{B_{R_{1}}^{c}}C(1+|x|^{2})\rho_{0}(x)\,\mathrm{d}x\leq\frac{7}{3}\eta+CR_{1}^{d}\exp(-\frac{1}{2}R_{1}^{2})<\frac{8}{3}\eta.$
Next, we apply the doubling of variable method [22, Theorem 8.3] and define a barrier function $\varphi:I_{t}\times\mathbb{R}^{d}\times I_{t}\times\mathbb{R}^{d}\to\mathbb{R}$
@@
Here, $\gamma,\varepsilon,\delta,\lambda\in(0,1)$ are parameters and we will specify their value later (cf. Figure 9). $l>2$ is a constant. The reader may assume $l=4$ for simplicity. We present the proof for a general $l$ so that the argument could be extended to general growth condition for the value function. $|x|_{l}^{l}:=\sum_{i=1}^{d}|x_{i}|^{l}$ denotes the $l$-Euclidean norm. We observe that all the terms in (B.20) are nonnegative.
Then, we define $\Phi_{k}:I_{t}\times\mathbb{R}^{d}\times I_{t}\times\mathbb{R}^{d}\to\mathbb{R}$
$\Phi_{k}(t,x,s,y):=V_{k}^{\iota}(t,x)-V_{k,\iota}^{*}(s,y)-\varphi(t,x,s,y)$ (B.21)
where for any $(t,x)\in[0,T]\times\mathbb{R}^{d}$
@@
$V_{k,\iota}^{*}(t,x)$ $:=\inf_{(s,y)\in[0,T]\times\mathbb{R}^{d}}\Big[V_{k}(s,y)+\frac{1}{2\iota}\left(|t-s|^{2}+|x-y|^{2}\right)\Big]$
are the Moreau envelopes for $V_{k}$ and $V_{k}^{*}$ respectively. The value of $\iota$ will be specified later (cf. Figure 9). Since $l>2$ and the value function grows at mostly quadratically in $|x|$ or $|y|$,
$\lim_{|x|\vee|y|\to\infty}\Phi_{k}(t,x,s,y)=-\infty.$
Therefore, $\Phi_{k}$ achieves a maximum $(t_{k},x_{k},s_{k},y_{k})\in I_{t}\times\mathbb{R}^{d}\times I_{t}\times\mathbb{R}^{d}$.
Since $V_{k}^{\iota}$ is semiconvex and $V_{k,\iota}^{*}$ is semiconcave, the function $V_{k}^{\iota}(t,x)-V_{k,\iota}^{*}(s,y)$ is semiconvex in $(t,x,s,y)$. Therefore, $V_{k}^{\iota}$ and $V_{k,\iota}^{*}$ are twice differentiable almost everywhere [2]. We remark that this differentiability is the reason to apply the Moreau envelope to the value function. In contrast, standard Schauder estimate (see the next step) only gives $C^{1}$ Hölder continuity for the value function in $t$.
Step 2. We present some estimations on $(t_{k},x_{k},s_{k},y_{k})$ in this step. Since $\Phi_{k}(t_{k},x_{k},s_{k},y_{k})\geq\Phi_{k}(t_{+},0,t_{+},0)$,
@@
We set $\lambda\leq\Delta t$ for $\lambda$ (cf. Figure 9). Then, $\gamma\Delta t\left(|x_{k}|_{l}^{l}+|y_{k}|_{l}^{l}\right)\leq C\left(1+|x_{k}|^{2}+|y_{k}|^{2}\right)$, which implies
@@
We denote $R_{2}:=C_{0}(\gamma\Delta t)^{-\frac{1}{l-2}}+2$, then $|x_{k}|,|y_{k}|\leq R_{2}-2$. Substitute (B.23) back into (B.22), we obtain
@@
where we record $c_{1}$ to decide the parameters later.
Next, we know from $2\Phi_{k}(t_{k},x_{k},s_{k},y_{k})\geq\Phi_{k}(t_{k},x_{k},t_{k},x_{k})+\Phi_{k}(s_{k},y_{k},s_{k},y_{k})$ that
@@
Therefore
@@
This implies,
$|t_{k}-s_{k}|+|x_{k}-y_{k}|\leq C_{1}(\gamma\Delta t)^{-\frac{2}{l-2}}(\varepsilon+\delta)$ (B.25)
(where we record this constant $C_{1}$ to decide $\varepsilon$, $\delta$ later) and
$\frac{1}{\varepsilon}|t_{k}-s_{k}|^{2}+\frac{1}{\delta}|x_{k}-y_{k}|^{2}\leq C(\gamma\Delta t)^{-\frac{4}{l-2}}(\varepsilon+\delta).$ (B.26)
Within a compact range $I_{t}\times B_{R_{2}}\times I_{t}\times B_{R_{2}}$, the term $\gamma(t_{+}-t+\Delta t)|x|_{l}^{l}+\gamma(t_{+}-s+\Delta t)|y|_{l}^{l}$ in (B.20) is semiconcave, so $\varphi$ is semiconcave. As a result, $\Phi_{k}$ is semiconvex in $I_{t}\times B_{R_{2}}\times I_{t}\times B_{R_{2}}$.
Next, we present a Schauder estimation for the value functions $V_{k}$ and $V_{k}^{*}$ within the compact domain $0,T]\times B_{R_{2}}$. We fix $\zeta\in(0,1)$ as the Hölder constant. For a function $V:[0,T]\times B_{R_{2}}\to\mathbb{R}$, we review the notation for Hölder semi-norm and Hölder norm (see [38, Chapter 4] for details):
$V]_{\zeta/2,\zeta}:=\sup_{(t,x)\neq(s,y)}\dfrac{|V(t,x)-V(s,y)|}{(|t-s|^{\frac{1}{2}}+|x-y|)^{\zeta}},\quad[V]_{1+\zeta/2,2+\zeta}:=[\partial_{t}V]_{\zeta/2,\zeta}+\sum_{i,j=1}^{d}[\partial_{x_{i}}\partial_{x_{j}}V]_{\zeta/2,\zeta},$
@@
We denote $f_{k}(t,x):=f(t,x,\mu^{k}_{t},\alpha_{k}(t,x))$ and define $b_{k}$ and $\sigma_{k}$ similarly. Then, for any $t,s\in[0,T]$, $x,y\in B_{R_{2}}$, we have
@@
$\leq K\left(1+R_{2}^{2}+W_{2}(\mu^{k}_{s},\delta_{0})^{2}+|\alpha_{k}(s,y)|^{2}\right)|t-s|^{\frac{1}{2}}+K(1+R_{2}+W_{2}(\mu^{k}_{t},\delta_{0})\vee W_{2}(\mu^{k}_{s},\delta_{0})$
$\quad+|\alpha_{k}(t,x)|\vee|\alpha_{k}(s,y)|)\left(|x-y|+W_{2}(\mu^{k}_{t},\mu^{k}_{s})+|\alpha_{k}(t,x)-\alpha_{k}(s,y)|\right)$
$\leq C(1+R_{2}^{2})|t-s|^{\frac{1}{2}}+C(1+R_{2})(|x-y|+|t-s|^{\frac{1}{2}})$
$\leq C_{R_{2}}\left(|t-s|^{\frac{1}{2}}+|x-y|\right),$
where $C_{R_{2}}>0$ is a constant that depend on $R_{2}$. Since $B_{R_{2}}$ is bounded, this implies $\left\|f_{k}\right\|_{\zeta/2,\zeta}\leq C_{R_{2}}$ within $0,T]\times B_{R_{2}}$. Similar estimations also hold for $b_{k}$ and $\sigma_{k}$. Therefore, applying standard Hölder estimation for linear parabolic equations [38, Section 4.5], we have
@@
Additionally, apply the Schauder estimation for the HJB equation [45], we have
@@
Step 3. We show that $t_{k}$, $s_{k}$ are not close to $t_{+}$ and $t_{-}$ in this step. Since $\Phi_{k}(\bar{t},\bar{x}_{k},\bar{t},\bar{x}_{k})\leq\Phi_{k}(t_{k},x_{k},s_{k},y_{k})$, we have
@@
$V_{k}^{\iota}(\bar{t},\bar{x}_{k})-V_{k,\iota}^{*}(\bar{t},\bar{x}_{k})\geq\frac{7}{3}\eta$ (B.30)
Then, we set $\gamma$ and $\lambda$ small enough s.t. (cf. Figure 9)
@@
Substitute (B.30) and (B.31) into (B.29), we get
$\quad\frac{5}{3}\eta\leq\Phi_{k}(t_{k},x_{k},s_{k},y_{k})=V_{k}^{\iota}(t_{k},x_{k})-V_{k,\iota}^{*}(s_{k},y_{k})-\varphi(t_{k},x_{k},s_{k},y_{k})$ (B.32)
$\leq V_{k}(t_{k},x_{k})-V_{k}^{*}(s_{k},y_{k})+C\iota R_{2}^{2}(2R_{2}^{2}+1)-\gamma\Delta t\left(|x_{k}|^{l}_{l}+|y_{k}|^{l}_{l}\right)-\frac{1}{2\varepsilon}|t_{k}-s_{k}|^{2}-\frac{1}{2\delta}|x_{k}-y_{k}|^{2}$
$\leq V_{k}(t_{+},x_{k})-V_{k}^{*}(t_{+},y_{k})+C(t_{+}-t_{k})(1+|x_{k}|^{2})+C(t_{+}-s_{k})(1+|y_{k}|^{2})$
$\quad+C_{2}\iota R_{2}^{2}(2R_{2}^{2}+1)-\frac{1}{2\delta}|x_{k}-y_{k}|^{2},$
where we used (A.49) and $t_{+}-t_{k}+\Delta t,t_{+}-s_{k}+\Delta t\geq\Delta t$ in the second inequality, used Lemma A.12 in the third inequality. Here, we record the constant $C_{2}$ in the last line. We set $\iota$ to be small enough (cf. Figure 9) s.t.
$C_{2}\iota R_{2}^{2}(2R_{2}^{2}+1)\leq\frac{1}{3}\eta.$ (B.33)
For the term $V_{k}(t_{+},x_{k})-V_{k}^{*}(t_{+},y_{k})$ in (B.32), we claim that up to a subsequence, we have
$V_{k}(t_{+},y_{k})-V_{k}^{*}(t_{+},y_{k})\leq\frac{1}{3}\eta\quad\forall k\geq 1.$ (B.34)
Otherwise, we have $\limsup_{k\to\infty}V_{k}(t_{+},y_{k})-V_{k}^{*}(t_{+},y_{k})\geq\frac{1}{3}\eta$, and we can subtract a subsequence (without changing the index $k$) s.t.
$V_{k}(t_{+},y_{k})-V_{k}^{*}(t_{+},y_{k})\geq\frac{1}{4}\eta\quad\forall k\geq 1.$ (B.35)
Let $r_{1}=\min\{\eta/(16CR_{2}),1\}$ where this $C$ is the constant in the inequality $|\nabla_{x}V(t,x)|\leq C(1+|x|)$ in Lemma A.12. Since $|y_{k}|\leq R_{2}-2$, for any $x$ s.t. $|x-y_{k}|\leq r_{1}$, we have $|x|\leq R_{2}-1$ and
@@
Substitute this into (B.35), we get
$V_{k}(t_{+},x)-V_{k}^{*}(t_{+},x)\geq\frac{1}{8}\eta\quad\forall k\geq 1,\,\,\forall x~\text{s.t.}~|x-y_{k}|\leq r_{1}.$
Taking integration, we get
@@
$\geq\int_{|x-y_{k}|\leq r_{1}}\frac{1}{8}\eta\,(2\pi)^{-d/2}\exp(-R_{2}^{2}/2)\,\mathrm{d}x=\left|B_{r_{1}}\right|\frac{1}{8}\eta\,(2\pi)^{-d/2}\exp(-R_{2}^{2}/2)>0,$
which contradicts to $\lim_{k\to\infty}h_{k}(t_{+})=0$. Therefore, the claim holds. We will take this subsequence s.t. (B.34) holds without changing the index $k$. We substitute (B.33) and (B.34) into (B.32), we get
@@
$\leq\frac{1}{2}C^{2}R_{2}^{2}\delta+CR_{2}^{2}[(t_{+}-t_{k})+(t_{+}-s_{k})]$
where we used Lemma A.12 and $|x_{k}|,|y_{k}|\leq R_{2}-2$ in the second inequality. We set $\delta$ to be small enough (cf. Figure 9) s.t. the first term in the last line of (B.36) satisfies $\frac{1}{2}C^{2}R_{2}^{2}\delta\leq\frac{1}{3}\eta$. Then
$\frac{2}{3}\eta\leq CR_{2}^{2}[(t_{+}-t_{k})+(t_{+}-s_{k})]~~\Rightarrow~~(t_{+}-t_{k})\vee(t_{+}-s_{k})\geq c_{2}\eta/R_{2}^{2},$ (B.37)
where we record this constant $c_{2}$ to decide other parameters. Recall we showed $|t_{k}-s_{k}|\leq C_{1}(\gamma\Delta t)^{-\frac{2}{l-2}}(\varepsilon+\delta)$ in (B.25). We set $\varepsilon$, $\delta$ small enough s.t. $\varepsilon,\delta\leq\frac{1}{4}C_{1}^{-1}(\gamma\Delta t)^{\frac{2}{l-2}}c_{2}\eta/R_{2}^{2}$. Then, we have $|t_{k}-s_{k}|\leq c_{2}\eta/(2R_{2}^{2})$ and hence
$(t_{+}-t_{k}),(t_{+}-s_{k})\geq c_{2}\eta/(2R_{2}^{2}).$ (B.38)
Recall that we also showed $(t_{k}-t_{-}),(s_{k}-t_{-})\geq c_{1}\lambda(\gamma\Delta t)^{\frac{2}{l-2}}$ in (B.24). Therefore, $(t_{k},x_{k},s_{k},y_{k})$ lies the interior of $I_{t}\times B_{R_{2}}\times I_{t}\times B_{R_{2}}$. We denote $r_{2}:=\frac{1}{2}\min\left\{c_{1}\lambda(\gamma\Delta t)^{\frac{2}{l-2}},c_{2}\eta/(2R_{2}^{2}),1\right\}$, then
@@
Step 4. In this step, we make a perturbation to the system. Let $\mu>0$, whose value will be specified later. Then the map
$(t,x,s,y)\mapsto\Phi_{k}(t,x,s,y)-\frac{\mu}{2}|(t,x,s,y)-(t_{k},x_{k},s_{k},y_{k})|^{2}$
achieves a strict maximum at $(t_{k},x_{k},s_{k},y_{k})$. Then, we make a small linear perturbation to the map above. Let $(q,p,\widehat{q},\widehat{p})\in\mathbb{R}^{1+d+1+d}$, we define
@@
We denote the maximum for $\widehat{\Phi}_{k}$ as $(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})$. Then, $(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})$ must lie in the set
[…]
which implies
@@
$\left|(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})-(t_{k},x_{k},s_{k},y_{k})\right|\leq r_{3}$ (B.40)
where $r_{3}>0$ will be specified later. We first require $2r_{3}\leq r_{2}$ (cf. Figure 9), which guarantees that
$\widehat{t}_{k},\widehat{s}_{k}\in[t_{-}+1.5r_{2},t_{+}-1.5r_{2}]$ (B.41)
because of (B.39). This also guarantees that $|\widehat{x}_{k}|,|\widehat{y}_{k}|\leq R_{2}-1$. By the optimality of $(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})$, we have
@@
where the gradient $\nabla$ is taken with respect to $(t,x,s,y)$. The critical point equation (B.42) shall be viewed as the expression for (q,p,q^,p^)($\widehat{q},\widehat{p}$) in terms of $(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})$. In order to bound (q,p,q^,p^)($\widehat{q},\widehat{p}$) in terms of $r_{3}$ (cf. (B.40)), we estimate $\left|\nabla\Phi_{k}(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})\right|$ next. Since $\nabla\Phi_{k}(t_{k},x_{k},s_{k},y_{k})=0$, we have

### v2

@@ B.2 Convergence of the gap for value function
$\leq\frac{7}{3}\eta+\int_{B_{R_{1}}^{c}}C(1+|x|^{2})\rho_{0}(x)\,\mathrm{d}x\to\frac{7}{3}\eta<\frac{8}{3}\eta\quad\text{as}\quad R_{1}\to\infty,$
which contradicts the property of the further subsequence stated above.
Next, we apply the doubling of variable method [22, Theorem 8.3] and define a barrier function $\varphi:I_{t}\times\mathbb{R}^{d}\times I_{t}\times\mathbb{R}^{d}\ni(t,x,s,y)\to\varphi(t,x,s,y)\in\mathbb{R}$ as follows:
@@
Here, $\gamma,\varepsilon,\delta,\lambda\in(0,1)$ are parameters, whose values will be later specified (cf. Figure 9). $|x|_{l}^{l}:=\sum_{i=1}^{d}|x_{i}|^{l}$ denotes the $l$-Euclidean norm and $l>2$ is a constant. In the proof, we assume that $l$ is an even integer, e.g., $l=4$, for simplicity. Nevertheless, the proof remains valid for a general value of $l$ and the argument can be extended to general growth conditions of value functions.
We define a sequence of functions $\Phi_{k}:I_{t}\times\mathbb{R}^{d}\times I_{t}\times\mathbb{R}^{d}\to\mathbb{R}$ as follows:
$\Phi_{k}(t,x,s,y):=V_{k}^{\iota}(t,x)-V_{k,\iota}^{*}(s,y)-\varphi(t,x,s,y),$ (B.21)
where for any $(t,x)\in[0,T]\times\mathbb{R}^{d}$,
@@
$V_{k,\iota}^{*}(t,x)$ $:=\inf_{(s,y)\in[0,T]\times\mathbb{R}^{d}}\Big[V_{k}^{*}(s,y)+\frac{1}{2\iota}\left(|t-s|^{2}+|x-y|^{2}\right)\Big],$
denote the Moreau envelopes for $V_{k}$ and $V_{k}^{*}$ respectively, with the value of $\iota$ to be later specified (cf. Figure 9). Since $l>2$ and the value functions have quadratic growth, $\lim_{|x|\vee|y|\to\infty}\Phi_{k}(t,x,s,y)=-\infty,\ \forall k\geq 1.$ Therefore, $\Phi_{k}$ attains its maximum at some point $(t_{k},x_{k},s_{k},y_{k})\in I_{t}\times\mathbb{R}^{d}\times I_{t}\times\mathbb{R}^{d}$.
Since $V_{k}^{\iota}$ is semiconvex and $V_{k,\iota}^{*}$ is semiconcave (cf. Lemma A.16), the function $V_{k}^{\iota}(t,x)-V_{k,\iota}^{*}(s,y)$ is semiconvex in $(t,x,s,y)$. Therefore, $V_{k}^{\iota}$ and $V_{k,\iota}^{*}$ are twice differentiable almost everywhere [2]. We remark that, differentiability in time is the main reason why we are using the Moreau envelope of the value function. In contrast, standard Schauder estimate (see the next step) only provides $C^{1}$ Hölder continuity in time.
Step 2. We present estimations for $(t_{k},x_{k},s_{k},y_{k})$. Since $\Phi_{k}(t_{k},x_{k},s_{k},y_{k})\geq\Phi_{k}(t_{+},0,t_{+},0)$,
@@
Take $\lambda$ such that $\lambda\leq\Delta t$ (cf. Figure 9). Since $\gamma\Delta t\left(|x_{k}|_{l}^{l}+|y_{k}|_{l}^{l}\right)\leq C\left(1+|x_{k}|^{2}+|y_{k}|^{2}\right)$,
@@
Denote $R_{2}:=C_{0}(\gamma\Delta t)^{-\frac{1}{l-2}}+2$ so that $|x_{k}|,|y_{k}|\leq R_{2}-2$. Substituting (B.23) back into (B.22) yields
@@
where we record the constant $c_{1}$ for the specification of the parameters.
From $2\Phi_{k}(t_{k},x_{k},s_{k},y_{k})\geq\Phi_{k}(t_{k},x_{k},t_{k},x_{k})+\Phi_{k}(s_{k},y_{k},s_{k},y_{k})$, we conclude that
@@
Using the local Lipschitz property of the Moreau envelope (Lemma A.16),
@@
This implies
$|t_{k}-s_{k}|+|x_{k}-y_{k}|\leq C_{1}(\gamma\Delta t)^{-\frac{2}{l-2}}(\varepsilon+\delta),$ (B.25)
where we record the constant $C_{1}$ for later specification of $\varepsilon$ and $\delta$. In addition, $\frac{1}{\varepsilon}|t_{k}-s_{k}|^{2}+\frac{1}{\delta}|x_{k}-y_{k}|^{2}\leq C(\gamma\Delta t)^{-\frac{4}{l-2}}(\varepsilon+\delta)$.
Next, we present a Schauder estimation for the value functions $V_{k}$ and $V_{k}^{*}$ within the compact domain $0,T]\times B_{R_{2}}$. Denote by $\zeta\in(0,1)$ the Hölder constant. For a function $V:[0,T]\times B_{R_{2}}\to\mathbb{R}$, the parabolic Hölder semi-norm and Hölder norm (see [38, Chapter 4] for details) are defined as follows:
$V]_{\zeta/2,\zeta}:=\sup_{(t,x)\neq(s,y)}\dfrac{|V(t,x)-V(s,y)|}{(|t-s|^{\frac{1}{2}}+|x-y|)^{\zeta}},\quad[V]_{1+\zeta/2,2+\zeta}:=[\partial_{t}V]_{\zeta/2,\zeta}+\sum_{i=1}^{d}[\partial_{x_{i}}V]_{\zeta/2,\zeta}+\sum_{i,j=1}^{d}[\partial_{x_{i}}\partial_{x_{j}}V]_{\zeta/2,\zeta},$
@@
Denote $f_{k}(t,x):=f(t,x,\mu^{k}_{t},\alpha_{k}(t,x))$ and define $b_{k}$, $\sigma_{k}$ similarly. For any $t,s\in[0,T]$, $x,y\in B_{R_{2}}$, by Assumption 4.1 and Assumption 4.3,
@@
$\leq K\left(1+R_{2}^{2}+W_{2}(\mu^{k}_{t},\delta_{0})^{2}\vee W_{2}(\mu^{k}_{s},\delta_{0})^{2}+|\alpha_{k}(s,y)|^{2}\vee|\alpha_{k}(t,x)|^{2}\right)|t-s|^{\frac{1}{2}}+K(1+R_{2}+$
$\quad W_{2}(\mu^{k}_{t},\delta_{0})\vee W_{2}(\mu^{k}_{s},\delta_{0})+|\alpha_{k}(t,x)|\vee|\alpha_{k}(s,y)|)\left(|x-y|+W_{2}(\mu^{k}_{t},\mu^{k}_{s})+|\alpha_{k}(t,x)-\alpha_{k}(s,y)|\right)$
$\leq C(1+R_{2}^{2})|t-s|^{\frac{1}{2}}+C(1+R_{2})(|x-y|+|t-s|^{\frac{1}{2}})\leq C_{R_{2}}\left(|t-s|^{\frac{1}{2}}+|x-y|\right),$
where $C_{R_{2}}>0$ is a constant that depends on $R_{2}$. Since $B_{R_{2}}$ is bounded, this implies $f_{k}]_{\zeta/2,\zeta}\leq C_{R_{2}}$ within $0,T]\times B_{R_{2}}$. Similar estimations also hold for $b_{k}$ and $\sigma_{k}$. Therefore, applying standard Hölder estimation for linear parabolic equations [38, Section 4.5], we have
@@
Additionally, applying the Schauder estimation for the HJB equation [45] yields
@@
Step 3. We show that $t_{k}$, $s_{k}$ are not close to $t_{+}$ and $t_{-}$. Since $\Phi_{k}(\bar{t},\bar{x}_{k},\bar{t},\bar{x}_{k})\leq\Phi_{k}(t_{k},x_{k},s_{k},y_{k})$,
@@
$V_{k}^{\iota}(\bar{t},\bar{x}_{k})-V_{k,\iota}^{*}(\bar{t},\bar{x}_{k})\geq\frac{7}{3}\eta,\ \forall k\geq 1.$ (B.29)
We set $\gamma$ and $\lambda$ small enough such that (cf. Figure 9)
@@
Substituting (B.29) and (B.30) into (B.28) yields
$\frac{5}{3}\eta\leq\Phi_{k}(t_{k},x_{k},s_{k},y_{k})=V_{k}^{\iota}(t_{k},x_{k})-V_{k,\iota}^{*}(s_{k},y_{k})-\varphi(t_{k},x_{k},s_{k},y_{k})$ (B.31)
$\hskip-5.0pt\leq V_{k}(t_{k},x_{k})-V_{k}^{*}(s_{k},y_{k})+C\,\iota R_{2}^{4}-\gamma\Delta t\left(|x_{k}|^{l}_{l}+|y_{k}|^{l}_{l}\right)-\frac{1}{2\varepsilon}|t_{k}-s_{k}|^{2}-\frac{1}{2\delta}|x_{k}-y_{k}|^{2}$
$\hskip-5.0pt\leq V_{k}(t_{+},x_{k})-V_{k}^{*}(t_{+},y_{k})+C(t_{+}-t_{k})(1+|x_{k}|^{2})+C(t_{+}-s_{k})(1+|y_{k}|^{2})+C_{2}\,\iota R_{2}^{4}-\frac{1}{2\delta}|x_{k}-y_{k}|^{2},$
where the second inequality follows from (A.49) and the last inequality follows from Lemma A.12. Here, we record the constant $C_{2}$ and set $\iota$ to be small enough (cf. Figure 9) such that
$C_{2}\,\iota R_{2}^{4}\leq\frac{1}{3}\eta.$ (B.32)
For the sequence $\{V_{k}(t_{+},x_{k})-V_{k}^{*}(t_{+},y_{k})\}_{k=1}^{\infty}$ that appears in (B.31), we claim that: there exists a subsequence (which is still denoted by index $k$) such that
$V_{k}(t_{+},y_{k})-V_{k}^{*}(t_{+},y_{k})\leq\frac{1}{3}\eta,\ \forall k\geq 1.$ (B.33)
We prove this argument by contradiction. If the claim does not hold, $\limsup_{k\to\infty}V_{k}(t_{+},y_{k})-V_{k}^{*}(t_{+},y_{k})\geq\frac{1}{3}\eta$ so that we can extract a subsequence (which is still denoted by index $k$) such that
$V_{k}(t_{+},y_{k})-V_{k}^{*}(t_{+},y_{k})\geq\frac{1}{4}\eta,\ \forall k\geq 1.$ (B.34)
Let $r_{1}:=\min\{\eta/(16CR_{2}),1\}$, where $C$ is the constant in $|\nabla_{x}V(t,x)|\leq C(1+|x|)$ within Lemma A.12. Since $|y_{k}|\leq R_{2}-2$, for any $x$ such that $|x-y_{k}|\leq r_{1}$, we have $|x|\leq R_{2}-1$ and
@@
Substituting this into (B.34) yields: for any $x$ such that $|x-y_{k}|\leq r_{1}$, $V_{k}(t_{+},x)-V_{k}^{*}(t_{+},x)\geq\frac{1}{8}\eta,\ \forall k\geq 1.$ Integrating both sides yields
@@
$\geq\frac{1}{8}\eta\int_{|x-y_{k}|\leq r_{1}}\,(2\pi)^{-d/2}\exp(-R_{2}^{2}/2)\,\mathrm{d}x=\left|B_{r_{1}}\right|\frac{1}{8}\eta\,(2\pi)^{-d/2}\exp(-R_{2}^{2}/2)>0,$
which contradicts $h(t_{+})=\lim_{k\to\infty}h_{k}(t_{+})=0$. Therefore, the claim (B.33) is true.
In the following context, we take the subsequence such that (B.33) holds, while maintaining the notation of the index as $k$. Substituting (B.32) and (B.33) into (B.31) yields
@@
$\leq\frac{1}{2}C^{2}R_{2}^{2}\delta+CR_{2}^{2}[(t_{+}-t_{k})+(t_{+}-s_{k})],$
where we used Lemma A.12 and $|x_{k}|,|y_{k}|\leq R_{2}-2$. Set $\delta$ to be small enough (cf. Figure 9) such that $\frac{1}{2}C^{2}R_{2}^{2}\delta\leq\frac{1}{3}\eta$. Then
$\frac{2}{3}\eta\leq CR_{2}^{2}[(t_{+}-t_{k})+(t_{+}-s_{k})]~~\Rightarrow~~(t_{+}-t_{k})\vee(t_{+}-s_{k})\geq\frac{c_{2}\eta}{R_{2}^{2}}.$ (B.36)
We record the constant $c_{2}$ for specification of parameters. Recall from (B.25) that $|t_{k}-s_{k}|\leq C_{1}(\gamma\Delta t)^{-\frac{2}{l-2}}(\varepsilon+\delta)$. By setting $\varepsilon$, $\delta$ small enough such that $\varepsilon,\delta\leq\frac{1}{4}C_{1}^{-1}(\gamma\Delta t)^{\frac{2}{l-2}}c_{2}\eta/R_{2}^{2}$, we get $|t_{k}-s_{k}|\leq(c_{2}\eta)/(2R_{2}^{2})$ and $(t_{+}-t_{k}),(t_{+}-s_{k})\geq(c_{2}\eta)/(2R_{2}^{2})$. Recall from (B.24) that $(t_{k}-t_{-}),(s_{k}-t_{-})\geq c_{1}\lambda(\gamma\Delta t)^{\frac{2}{l-2}}$. Therefore, $(t_{k},x_{k},s_{k},y_{k})$ is an interior point of $I_{t}\times B_{R_{2}}\times I_{t}\times B_{R_{2}}$. By denoting $r_{2}:=\frac{1}{2}\min\left\{c_{1}\lambda(\gamma\Delta t)^{\frac{2}{l-2}},c_{2}\eta/(2R_{2}^{2}),1\right\}$,
@@
Step 4. In this step, we introduce perturbation to the system. Let $\mu>0$, whose value will be later specified. The mapping $(t,x,s,y)\mapsto\Phi_{k}(t,x,s,y)-\frac{\mu}{2}|(t,x,s,y)-(t_{k},x_{k},s_{k},y_{k})|^{2}$ attains a strict maximum at $(t_{k},x_{k},s_{k},y_{k})$. For $(q,p,\widehat{q},\widehat{p})\in\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{d}$, define
@@
whose maximum is attained by $(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})$. Then, $(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})$ must lie in the set
@@
$\left|(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})-(t_{k},x_{k},s_{k},y_{k})\right|\leq r_{3},$ (B.38)
where $0<2r_{3}\leq r_{2}$ (cf. Figure 9) will be later specified. Due to (B.37), this guarantees
$\widehat{t}_{k},\widehat{s}_{k}\in[t_{-}+1.5r_{2},t_{+}-1.5r_{2}].$ (B.39)
Recall that $r_{2}\leq\frac{1}{2}$, which implies $r_{3}\leq 1$ and $|\widehat{x}_{k}|,|\widehat{y}_{k}|\leq R_{2}-1$. The optimality of $(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})$ provides
[…]
where the gradient $\nabla$ is taken with respect to $(t,x,s,y)$. The critical point equation (B.40) expresses (q,p,q^,p^)($\widehat{q},\widehat{p}$) in terms of $(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})$. In order to bound |(q,p,q^,p^)||($\widehat{q},\widehat{p}$)| in terms of $r_{3}$ (cf. (B.38)), our next task is to estimate $\left|\nabla\Phi_{k}(\widehat{t}_{k},\widehat{x}_{k},\widehat{s}_{k},\widehat{y}_{k})\right|$. Using $\nabla\Phi_{k}(t_{k},x_{k},s_{k},y_{k})=0$,